OpenStudy (anonymous):

this isn't a question I just want to test something... is there something wrong with this equation? 5+5+5+5 = 555 think about it.

OpenStudy (compassionate):

i. It's an expression, not an equation ii. 5 + 5 + 5 + 5 = 20; not 555. You might've meant 5555

OpenStudy (anonymous):

no i mean 5+5+5+5 \[\neq555\] lol :p

OpenStudy (compassionate):

Of course 5 + 5 + 5 + 5 =/= 555. =P

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

No they're isn't an issue with the equation 5+5+5+5=20 and it's not 555